# Reinforcing the Diffusion Chain of Lateral Thought with Diffusion Language Models

### v1

@@ 1 Introduction
We conduct experiments on both Math and code generation tasks to demonstrate the efficacy of DCoLT. After training the DCoLT on the SEDD 400M model, it achieves 96.2% and 57.0% in accuracy on Sudoku 4$\times$4 and GSM8K-Aug [10] tasks. On the LLaDA 8B model, DCoLT achieves the state-of-the-art performance among existing DLMs trained with SFT or RL or even both. Using only public data and 16 H800 GPUs, the DCoLT-reinforced LLaDA model achieves 88.1% on GSM8K [8], 44.6% on MATH [16], 51.6% on MBPP [4] and 59.1% on HumanEval [7] in the challenging zero-shot setting. Even compared with autoregressive models that are trained with significantly more proprietary data and fully annotated CoT reasoning processes, it still demonstrates competitive performances.
@@ 2 Related Works
Among discrete diffusion models, masked diffusion models [27, 42, 24] emerge as a promising approach due to their superior performances. Recent advances simplify model training and design by adopting the straight cross-entropy loss [48, 29] and removing the timestep embeddings [47]. The resultant models [27, 42] achieve the state-of-the-art performance among existing diffusion language models.
@@
However, existing RL approaches are primarily applied to autoregressive language models [30, 15]. For DLMs, previous works explored policy gradient methods within the concrete score-matching framework [45], or estimated the log-probability of the sampled responses to perform reinforcement learning [46]. However, both models merely focused on the generated text response in the final diffusion step as an action to optimize, ignoring the role of intermediate diffusion steps as lateral thought to reinforce amid the reverse diffusion process.
@@ 3.1 Diffusion Chain of Lateral Thought
  • Nonlinear Generation: CoT generates tokens one-by-one linearly in an autoregressive manner from left to right. In contrast, DCoLT allows nonlinear generation of tokens at various positions. This aligns with how human develops ideas by beginning with keywords or critical points and then refining the details around them.
@@ 3.2 A Continuous-Time DLM Case: DCoLT-reinforced SEDD
Figure 2: The model structure of DCoLT-LLaDA. It first predicts the token set to unmask $\mathcal{U}_{n}$ according to the ranking score $h_{\theta,n}$, and then samples those unmasked tokens in $\mathcal{U}_{n}$.
@@ 3.3 A Discrete-Time DLM Case: DCoLT-reinforced LLaDA
Specifically, UPM takes the hidden states from the last hidden layer as inputs at each denoising step $n$, and predicts a ranking score $h_{\theta,n}^{i}$ for $i$-th token. It contains only one transformer block, introducing marginal computation cost. Moreover, the step index $n$ and the mask indicator of each token in $x_{n-1}$ can benefit the model training. Thus, we embed both of them into the Unmask Policy Module (UPM) via adaptive layer normalization. The module structure is shown in Fig. 2.
As shown in Fig. 10 of Appendix B, the learned scores $h_{\theta,n}$ can be viewed as the ranked confidences over the token predictions at the current step $n$. Higher scores indicate that the UPM predicts lower levels of diffusion noises may be present in the generated tokens, which are less likely to contain errors and thus could be unmasked with higher priorities at the current step. Figure 7 shows the resulting unmasking orders from the learned ranking scores for some prompts.
Once the unmask set $\mathcal{U}_{n}$ is determined, the model predicts their token values based on the output distribution by LLaDA over the vocabulary. Viewing this prediction as a second-stage action, the probability of generating $x_{n}$ given $x_{n-1}$ and $\mathcal{U}_{n}$ can be written as
@@ 4 Experiments
4.1 Experiments on DCoLT-SEDD
@@ 4.1.2 Results
4.2 Experiments on DCoLT-LLaDA
@@ 4.2.1 Settings
We further apply DCoLT to LLaDA [27], a state-of-the-art discrete-time masked-based DLM with 8B parameters. During generation, it starts with a masked sequence of length $256$, and performs a 256-step reverse diffusion process. We adopt a block-wise unmasking strategy as in LLaDA [27], and divide the sequence into blocks of length $8$. In Table 2, both LLaDA 8B and LLaDA-DCoLT 8B are evaluated with this setting for fair comparison. We conduct experiments with 16 H800 GPUs to jointly train both UPM and LLaDA parts. For implementation details, please refer to Appendix C.2.
@@ 4.2.2 Results
LLaDA + DCoLT 8B (Ours)    + RL 88.1 (0) 44.6 (0) 59.1 (0) 51.6 (0)
@@
As shown in Tab. 2, among all compared DLMs, DCoLT consistently achieves the best performance across all benchmarks. On GSM8K, DCoLT reaches 88.1%, significantly higher than other methods. On the more challenging MATH dataset, DCoLT achieves 44.6%, outperforming baseline models such as LLaDA 8B (+5.7%) and Dream 7B (+1.7%), as well as models enhanced with post-training by SFT, RL or both. Notably, d1-LLaDA [46] uses 1K questions paired with detailed reasoning traces [26] in SFT to boost reasoning capabilities. In contrast, DCoLT achieves superior performance – +6.0% on GSM8K and +4.4% on MATH – without any reasoning supervision, relying solely on reward signals based on the final answers.
Even when compared with autoregressive models trained with a significantly larger amount of proprietary data, DCoLT remains highly competitive. For example, DeepseekMath [36] gathers 776K questions with reasoning process annotations for SFT and 144K questions for RL, while DCoLT is trained on just 15K public training samples from GSM8k and MATH. Despite using nearly two orders of magnitude fewer data for RL training, DCoLT matches DeepseekMath’s performance on GSM8K (88.1%), highlighting its data efficiency and the effectiveness of reward-driven reasoning via diffusion-based lateral reasoning. In Appendix B, we analyze the thinking process generated by DCoLT in LLaDA 8B model.
Similar trends are observed on code generation benchmarks. Unlike models that rely on valuable ground-truth code for supervised training, DCoLT requires no code for supervised training at all. Instead, it only needs several test cases to provide a simple outcome-based reward for reinforcement learning: the model is rewarded only if the generated code passes all unit tests. Despite this code-free reward signal, DCoLT achieves 51.6% on MBPP and 59.1% on HumanEval, outperforming other DLMs and auto-regressive models, except for DeepseekCoder [14] that benefits from 2B tokens of high-quality instruction-tuning data—approximately 150×150$\times$150 × more than we used (about 12M tokens). Details of post-training datasets used by all compared methods are provided in Appendix C.3.
@@ 4.2.3 Ablation Studies
In this section, we provide some ablation studies to reveal the role of some model design and hyper-parameter settings in DCoLT. Unless otherwise specified, all experiments are conducted on LLaDA 8B with $N=64$ diffusion steps, and the model is trained for $150$ iterations with a batch size of $32$.
@@
In Table 3, we ablate the model design in DCoLT. Even if we freeze the original model parameters in LLaDA and only train the additional unmasking policy module, it achieves a significant improvement from 47.27% to 69.24%. This result indicates that the unmasking policy plays a crucial role in our model. By training the LLaDA part with RL together, the accuracy further improves to 81.06%.
@@
Though the models in Table 2 is trained to generate completions with a fixed generation length $256$ (i.e., the length of initialized mask tokens), it can benefit from generating longer sequences without further RL training on longer generations. This is especially effective when answering difficult questions on MATH dataset, in which solving problems sometimes require longer reasoning. Table 4 reports the accuracies with different generation lengths. It suggests that longer generations in inference lead to better accuracies. For example, a generation length of $384$ achieves an accuracy of 46.22%, which is +1.62% higher than the default length of $256$. This improvement vanishes as we further increase the generation length. However, if we further tune the model on a longer generation length, for example $L=512$, the accuracy further improves to 47.3%. This suggests the potential of how scaled generation lengths may improve the performance of DCoLT-trained DLMs, as shown in Figure 13 of Appendix D.
@@ Appendix A Analysis of Thinking Process in SEDD+DCoLT
In Fig 3, we plot the distribution of generation step for these three cell types, the left side displays statistics for the SEDD + SFT model, while the right side shows our further RL-trained SEDD + DCoLT model. The right figure reveals that the our SEDD + DCoLT significantly prioritizes generating easy cells in earlier steps than hard cells, as these cells can be directly determined from the initial given cells, indicating that our model learns a progressive generation strategy from easy to hard. Notably, in the left figure, we observe almost no difference between generation orders across various difficulty levels for SEDD + SFT model. This comparison reveals that the DCoLT changes its SFT-trained counterpart’s reasoning behavior to more progressive generation aligned naturally with how humankind handles problems in a real world.
@@
For GSM8K-Aug dataset, since it’s hard to determine which token position is more difficult to predict, we only show the generation process and the generation step statistics for each position. Fig. 5 demonstrates the contrasting generation behaviors between SEDD + DCoLT and GPT2 + CoT on the GSM8K-Aug dataset. While GPT2 + CoT consistently follows a left-to-right generation order (where earlier generation steps correspond to lower token positions), SEDD + DCoLT shows minimal positional bias in generation order across token positions. Furthermore, our results in Fig 6 reveal that SEDD + DCoLT’s generation order is sample-dependent, maintaining a nonlinear generation throughout the process.
@@
Appendix B Analysis of Thinking Process in LLaDA+DCoLT
We illustrate how LLaDA + DCoLT generates a 64-token answer to a GSM8K question in Fig. 7. Tokens shown in darker shades are generated in later diffusion steps. As the figure shows, key numbers and symbols tend to emerge early in the generation process, while surrounding textual elements are filled in later to ensure grammatical correctness and fluency.
@@ Appendix B Analysis of Thinking Process in LLaDA+DCoLT
Figure 7: Examples of 64-token responses generated by LLaDA+DCoLT on GSM8K. Token color reflects the generation order, progressing from light (early steps) to dark (later steps). The corresponding questions are provided in the sub-captions. We observe from the results that many tokens bearing key information to final answers such as the numbers, units and factual entities tend to be unmasked at earlier steps, while those tokens for function words and symbols such as “of" and “the" tend to be unmasked at later steps.
Figure 8: Intermediate step outputs from LLaDA+DCoLT during the generation of a 256-token response. The question is from GSM8K: A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?
Figure 9: Intermediate step outputs from LLaDA+DCoLT during the generation of a 256-token response. The question is from GSM8K: Cars have lined up on the motorway. Some of the cars drive through in the first 15 minutes of the traffic jam, then 20 more cars drive through in the remaining 15 minutes of the jam. 5 cars from the line take an exit so they don’t have to drive through the traffic jam. If there were originally 30 cars on the motorway, how many cars drove through the traffic jam in the first 15 minutes?
@@
Figure 10: Illustration of intermediate steps of generated responses by LLaDA+DCoLT on GSM8K. Gray tokens represent positions that have already been unmasked prior to the current step. For each masked position, we show the predicted token with the highest generation probability. Token color indicates the unmasking score, with darker blue indicating higher probability of be unmasked. The token eventually selected to unmask at each step is bounded in a red box. The corresponding prompts are given in the sub-captions.
@@ C.1 DCoLT-SEDD
We first generated 50,000 sudoku puzzles of size $4\times 4$ by uniformly replacing 1-9 cells with zeros. Using DeepSeek V3-0324[9], we produced CoT reasoning steps and filtered the results to remove repetitive responses and incorrect solutions. All training datasets were created from this cleaned dataset to ensure fair comparison. Dataset samples can be found in the Fig 11.
Figure 11: The CoT data samples, when apply to our DCoLT post-training, we only use the problems above without step-by-step process.
@@ C.2.2 Implementation Details for Code Generation
C.3 Post-Training Data Used in Table 2
In Table 7, we list the data used in post-training for math and code generation tasks by different models. Note that some methods [34, 14, 36] involve a considerable large amount of proprietary math/code data for both pretraining and post-training of models, which makes crucial contributions to high accuracies.
Table 7: Post-training data for math / code generation used in Table 2. “-" means this type of post-training was not used for the model.
@@ C.3 Post-Training Data Used in Table 2
LLaDA + DCoLT 8B (Ours) - Math: 7.5K GSM8K + 7.5K MATH / Code: 48K filtered from KodCode
@@ Appendix D Potential Scaling Law for Longer Generations
Figure 13 illustrates some examples of generated responses of various lengths to the same prompt for the DCoLT-reinforced LLaDA. In DLMs, the generation length tends to increase as more mask tokens are initialized at the beginning of the reverse diffusion process. This differs from autoregressive language models, where the length of a generated response is determined by when the End of Text (EoT) token appears.
Although DLMs also use the EoT token to mark the end of the output, the token often appears later in the output sequence to generate a longer thinking process when more mask tokens are initialized. This offers a flexible way for DLMs to control various lengths of generated responses. As shown in Table 4, longer generations could improve performance, particularly when the model is fine-tuned by DCoLT with increasing generation length. This suggests the existence of a potential scaling law for longer generations.
Figure 13(a) shows that some mistakes generated in the thinking process can be corrected in longer generations, thus achieving better results over shorter ones. We will study such a scaling phenomenon further in the future.

### v2

@@ 1 Introduction
We conduct experiments on both Math and code generation tasks to demonstrate the efficacy of DCoLT. After training the DCoLT on the SEDD 400M model, it achieves 96.2% and 57.0% in accuracy on Sudoku 4$\times$4 and GSM8K-Aug [10] tasks. On the LLaDA 8B model, DCoLT achieves the state-of-the-art performance among existing DLMs trained with SFT or RL or even both. Using only public data and 16 H800 GPUs, the DCoLT-reinforced LLaDA model achieves 88.1% on GSM8K [8], 44.6% on MATH [16], 51.6% on MBPP [4] and 59.1% on HumanEval [7] in the challenging zero-shot setting. Even compared with auto-regressive models that are trained with significantly more proprietary data and fully annotated CoT reasoning processes, it still demonstrates competitive performances.
@@ 2 Related Works
Among discrete diffusion models, masked diffusion models [27, 42, 24] emerge as a promising approach due to their superior performances. Recent advances simplify model training and design by adopting the straight cross-entropy loss [48, 29] and removing the time embeddings [47]. The resultant models [27, 42] achieve the state-of-the-art performance among existing diffusion language models.
@@
However, existing RL approaches are primarily applied to auto-regressive language models [30, 15]. For DLMs, previous works explored policy gradient methods within the concrete score-matching framework [45], or estimated the log-probability of the sampled responses to perform reinforcement learning [46]. However, both models merely focused on the generated text response in the final diffusion step as an action to optimize, ignoring the role of intermediate diffusion steps as lateral thought to reinforce amid the reverse diffusion process.
@@ 3.1 Diffusion Chain of Lateral Thought
  • Nonlinear Generation: CoT generates tokens one-by-one linearly in an auto-regressive manner from left to right. In contrast, DCoLT allows nonlinear generation of tokens at various positions. This aligns with how human develops ideas by beginning with keywords or critical points and then refining the details around them.
@@ 3.2 A Continuous-Time DLM Case: DCoLT-reinforced SEDD
Figure 2: The structure of LLaDOU. It first predicts the token set to unmask $\mathcal{U}_{n}$ according to the ranking score $h_{\theta,n}$ by the UPM, and then samples those unmasked tokens in $\mathcal{U}_{n}$ by LLaDA blocks.
@@ 3.3 A Discrete-Time DLM Case: DCoLT-reinforced LLaDA
Specifically, UPM takes the hidden states from the last hidden layer as inputs at each denoising step $n$, and predicts a ranking score $h_{\theta,n}^{i}$ for $i$-th token. It contains only one transformer block, introducing marginal computation cost. Moreover, the step index $n$ and the mask indicator of each token in $x_{n-1}$ can benefit the model training. Thus, we embed both of them into the Unmask Policy Module (UPM) via adaptive layer normalization. The module structure is shown in Fig. 2. For convenience, we name the DCoLT-trained LLaDA with UPM by LLaDOU (LLaDA with Ordered Unmasking).
As shown in Fig. 10 of Appendix B, the learned scores $h_{\theta,n}$ can be viewed as the ranked confidences over the token predictions at the current step $n$. Higher scores indicate that the UPM predicts lower levels of diffusion noises may be present in the generated tokens, which are less likely to contain errors and thus could be unmasked with higher priorities at the current step. Fig. 7 shows the resulting unmasking orders from the learned ranking scores for some prompts.
Once the unmask set $\mathcal{U}_{n}$ is determined, the model predicts their token values based on the output distribution by LLaDA blocks over the vocabulary. Viewing this prediction as a second-stage action, the probability of generating $x_{n}$ given $x_{n-1}$ and $\mathcal{U}_{n}$ can be written as
@@ 4 Experiments
4.1 Experiments on SEDD+DCoLT
@@ 4.1.2 Results
4.2 Experiments on LLaDOU (LLaDA+DCoLT)
@@ 4.2.1 Settings
We further apply DCoLT to LLaDA [27], a SOTA discrete-time masked-based DLM with 8B parameters. During generation, it starts with a masked sequence of length $256$, and performs a 256-step reverse diffusion process. We adopt a block-wise unmasking strategy as in LLaDA [27], and divide the sequence into blocks of length $8$. In Tab. 2, both LLaDA and LLaDOU (LLaDA+DCoLT) are evaluated with this setting for fair comparison. We conduct experiments with 16 H800 GPUs to jointly train both UPM and LLaDA parts. For implementation details, please refer to Appendix C.2.
@@ 4.2.2 Results
LLaDOU (LLaDA + DCoLT) 8B    + RL 88.1 (0) 44.6 (0) 59.1 (0) 51.6 (0)
@@
As shown in Tab. 2, among all compared DLMs, LLaDOU consistently achieves the best performance across all benchmarks. On GSM8K, LLaDOU reaches 88.1%, significantly higher than other methods. On the more challenging MATH dataset, LLaDOU achieves 44.6%, outperforming baseline models such as LLaDA 8B (+5.7%) and Dream 7B (+1.7%), as well as models enhanced with post-training by SFT, RL or both. Notably, d1-LLaDA [46] uses 1K questions paired with detailed reasoning traces [26] in SFT to boost reasoning capabilities. In contrast, LLaDOU achieves superior performance – +6.0% on GSM8K and +4.4% on MATH – without any reasoning supervision, relying solely on reward signals based on the final answers.
Even when compared with auto-regressive models trained with a significantly larger amount of proprietary data, LLaDOU remains highly competitive. For example, DeepseekMath [36] gathers 776K questions with reasoning process annotations for SFT and 144K questions for RL, while LLaDOU is trained on just 15K public training samples from GSM8K and MATH. Despite using nearly two orders of magnitude fewer data for RL training, LLaDOU matches DeepseekMath’s performance on GSM8K (88.1%), highlighting its data efficiency and the effectiveness of reward-driven reasoning via diffusion-based lateral reasoning. In Appendix B, we analyze the thinking process generated by LLaDOU.
Similar trends are observed on code generation benchmarks. Unlike models that rely on valuable ground-truth code for supervised training, LLaDOU requires no code for supervised training at all. Instead, it only needs several test cases to provide a simple outcome-based reward for reinforcement learning: the model is rewarded only if the generated code passes all unit tests. Despite this code-free reward signal, LLaDOU achieves 51.6% on MBPP and 59.1% on HumanEval, outperforming other DLMs and auto-regressive models, except for DeepseekCoder [14] that benefits from 2B tokens of high-quality instruction-tuning data—approximately 150×150$\times$150 × more than we used (about 12M tokens). Details of post-training datasets used by all compared methods are provided in Appendix C.3.
@@ 4.2.3 Ablation Studies
We provide some ablation studies to reveal the role of some model design and hyper-parameter settings in LLaDOU. Unless otherwise specified, all experiments are conducted on LLaDOU 8B with $N=64$ diffusion steps, and the model is trained for $150$ iterations with a batch size of $32$.
@@
In Tab. 3, we ablate the model design in LLaDOU. Even if we freeze the original model parameters in LLaDA part and only train the UPM part, it achieves a significant improvement from 47.27% to 69.24%. This result indicates that the unmasking policy plays a crucial role in our model. By training the LLaDA part with RL together, the accuracy further improves to 81.06%.
@@
Though the models in Tab. 2 is trained to generate completions with a fixed generation length $256$ (i.e., the length of initialized mask tokens), it can benefit from generating longer sequences without further RL training on longer generations. This is especially effective when answering difficult questions on MATH dataset, in which solving problems sometimes require longer reasoning. Tab. 4 reports the accuracies with different generation lengths. It suggests that longer generations in inference lead to better accuracies. For example, a generation length of $384$ achieves an accuracy of 46.22%, which is +1.62% higher than the default length of $256$. This improvement vanishes as we further increase the generation length. However, if we further tune the model on a longer generation length, for example $L=512$, the accuracy further improves to 47.3%. This suggests the potential of how scaled generation lengths may improve the performance of DCoLT-trained DLMs, as shown in Fig. 13 of Appendix D.
@@ Appendix A Analysis of Thinking Process in SEDD+DCoLT
In Fig. 3, we plot the distribution of generation step for these three cell types, the left side displays statistics for the SEDD + SFT model, while the right side shows our further RL-trained SEDD + DCoLT model. The right figure reveals that the our SEDD + DCoLT significantly prioritizes generating easy cells in earlier steps than hard cells, as these cells can be directly determined from the initial given cells, indicating that our model learns a progressive generation strategy from easy to hard. Notably, in the left figure, we observe almost no difference between generation orders across various difficulty levels for SEDD + SFT model. This comparison reveals that the DCoLT changes its SFT-trained counterpart’s reasoning behavior to more progressive generation aligned naturally with how humankind handles problems in a real world.
@@
For GSM8K-Aug dataset, since it’s hard to determine which token position is more difficult to predict, we only show the generation process and the generation step statistics for each position. Fig. 5 demonstrates the contrasting generation behaviors between SEDD + DCoLT and GPT2 + CoT on the GSM8K-Aug dataset. While GPT2 + CoT consistently follows a left-to-right generation order (where earlier generation steps correspond to lower token positions), SEDD + DCoLT shows minimal positional bias in generation order across token positions. Furthermore, our results in Fig. 6 reveal that SEDD + DCoLT’s generation order is sample-dependent, maintaining a nonlinear generation throughout the process.
@@
Appendix B Analysis of Thinking Process in LLaDOU (LLaDA+DCoLT)
We illustrate how LLaDOU generates a 64-token answer to a GSM8K question in Fig. 7. Tokens shown in darker shades are generated in later diffusion steps. As the figure shows, key numbers and symbols tend to emerge early in the generation process, while surrounding textual elements are filled in later to ensure grammatical correctness and fluency.
@@ Appendix B Analysis of Thinking Process in LLaDOU (LLaDA+DCoLT)
Figure 7: Examples of 64-token responses generated by LLaDOU on GSM8K. Token color reflects the generation order, progressing from light (early steps) to dark (later steps). The corresponding questions are provided in the sub-captions. We observe from the results that many tokens bearing key information to final answers such as the numbers, units and factual entities tend to be unmasked at earlier steps, while those tokens for function words and symbols such as “of" and “the" tend to be unmasked at later steps.
Figure 8: Intermediate step outputs from LLaDOU during the generation of a 256-token response. The question is from GSM8K: A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?
Figure 9: Intermediate step outputs from LLaDOU during the generation of a 256-token response. The question is from GSM8K: Cars have lined up on the motorway. Some of the cars drive through in the first 15 minutes of the traffic jam, then 20 more cars drive through in the remaining 15 minutes of the jam. 5 cars from the line take an exit so they don’t have to drive through the traffic jam. If there were originally 30 cars on the motorway, how many cars drove through the traffic jam in the first 15 minutes?
@@
Figure 10: Illustration of intermediate steps of generated responses by LLaDOU on GSM8K. Gray tokens represent positions that have already been unmasked prior to the current step. For each masked position, we show the predicted token with the highest generation probability. Token color indicates the unmasking score, with darker blue indicating higher probability of be unmasked. The token eventually selected to unmask at each step is bounded in a red box. The corresponding prompts are given in the sub-captions.
@@ C.1 SEDD+DCoLT
We first generated 50,000 sudoku puzzles of size $4\times 4$ by uniformly replacing 1-9 cells with zeros. Using DeepSeek V3-0324[9], we produced CoT reasoning steps and filtered the results to remove repetitive responses and incorrect solutions. All training datasets were created from this cleaned dataset to ensure fair comparison. Dataset samples can be found in the Fig. 11.
Figure 11: Examples of CoT data used for training GPT2+CoT and SEDD+DoT.
@@ C.2.2 Implementation Details for Code Generation
C.3 Post-Training Data Used in Tab. 2
In Tab. 7, we list the data used in post-training for math and code generation tasks by different models. Note that some methods [34, 14, 36] involve a considerable large amount of proprietary math/code data for both pretraining and post-training of models, which makes crucial contributions to high accuracies.
Table 7: Post-training data for math / code generation used in Tab. 2. “-" means this type of post-training was not used for the model.
@@
LLaDOU (LLaDA + DCoLT) 8B - Math: 7.5K GSM8K + 7.5K MATH / Code: 48K filtered from KodCode
@@ Appendix D Potential Scaling Law for Longer Generations
Fig. 13 illustrates some examples of generated responses of various lengths to the same prompt for the LLaDOU. In DLMs, the generation length tends to increase as more mask tokens are initialized at the beginning of the reverse diffusion process. This differs from auto-regressive language models, where the length of a generated response is determined by when the End of Text (EoT) token appears.
Although DLMs also use the EoT token to mark the end of the output, the token often appears later in the output sequence to generate a longer thinking process when more mask tokens are initialized. This offers a flexible way for DLMs to control various lengths of generated responses. As shown in Tab. 4, longer generations could improve performance, particularly when the model is fine-tuned by DCoLT with increasing generation length. This suggests the existence of a potential scaling law for longer generations.
Fig. 13(a) shows that some mistakes generated in the thinking process can be corrected in longer generations, thus achieving better results over shorter ones. We will study such a scaling phenomenon further in the future.